OpenStudy (anonymous):

Hello Guys .. Can You help me .. this is for my project in school . can you give me example of arithmetic progration and geometric progration with its formula ...... please help me because the deadline of our project is for tomorrow .. thanks :)) lol

OpenStudy (anonymous):

its progression right?

OpenStudy (anonymous):

yah ..

OpenStudy (anonymous):

no i mean its PROGRATION .. as in sequence !

OpenStudy (anonymous):

A sequence of numbers in which each differs from the preceding by a constant quantity (e.g., 3, 6, 9, 12, etc.; 9, 7, 5, 3, etc.). arithmetic progression

OpenStudy (anonymous):

A progression of numbers with a constant ratio between each number and the one before (e.g. 1, 3, 9, 27, 81) this is geometric progression

OpenStudy (anonymous):

give me an example of arith metic and geometric progression with its formula .. :))

OpenStudy (anonymous):

a geometric sequence can be written as: aq0=a, aq1=aq, aq2, q3, ... where q ≠ 0, q is the common ratio and a is a scale factor. Formulae for the n-th term can be defined as: an = an-1.q an = a1.qn-1 The common ratio then is: q = ak ak-1 A sequence with a common ratio of 2 and a scale factor of 1 is 1, 2, 4, 8, 16, 32... A sequence with a common ratio of -1 and a scale factor of 5 is 5, -5, 5, -5, 5, -5,... If the common ratio is: Negative, the results will alternate between positive and negative. Greater than 1, there will be exponential growth towards infinity (positive). Less than -1, there will be exponential growth towards infinity (positive and negative). Between 1 and -1, there will be exponential decay towards zero. Zero, the results will remain at zero Geometric Progression Properties a2k = ak-1.ak+1 a1.an = a2.an-1 =...= ak.an-k+1 Formula for the sum of the first n numbers of a geometric series Sn = a1 - anq 1 - q = a1. 1 - qn 1 - q Infinite geometric series where |q| < 1 If |q| < 1 then an -> 0, when n -> ∞. The sum S of such an infinite geometric series is: S = a1 1 1 - x which is valid only for |x| < 1 and a1 is the first term.

hartnn (hartnn):

Hello, @kimii , \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) you can find formulas and an example in this link , try it...if you still need more examples,just ask. http://openstudy.com/study#/updates/503bb2a0e4b007f9003103b0